OpenStudy (anonymous):

does anyone know how to solve x^2-k=0 if k>0 HELP!

OpenStudy (anonymous):

\[x^2-k=0\]add k to both sides\[x^2=k\]take square root

OpenStudy (anonymous):

it confuses me so??

OpenStudy (anonymous):

\[\sqrt{x^2}=\sqrt{k}\]

OpenStudy (anonymous):

\[|x|=\sqrt{k}\]since\[\sqrt{x^2}=|x|\]

OpenStudy (anonymous):

so what is the final answer? what do u think?

OpenStudy (anonymous):

x=?

OpenStudy (anonymous):

my choices are \[\sqrt{k}\]or k or \[\pm \sqrt{k}\]or -\[\sqrt{k}\]

OpenStudy (anonymous):

im still not sure

OpenStudy (anonymous):

I think it is sq.rtk

OpenStudy (anonymous):

am I right?